# Revenue Maximization Under Sequential Price Competition Via The Estimation Of $s$-Concave Demand Functions

### v3

@@ Appendix H NPLS for $s$-concave regression
Inspired by the work of Dümbgen et al. (2004), we prove a uniform convergence result for the NPLSE of a mean function that can be written as an increasing function $h$ composed with a concave function $\phi_{0}$. Suppose that we observe $\left(U_{1},Y_{1}\right),\left(U_{2},Y_{2}\right),\ldots,\left(U_{n},Y_{n}\right)$ with fixed numbers $U_{1}\leq U_{2}\leq$ $\cdots\leq U_{n}$ and independent random variables $Y_{1},Y_{2},\ldots,Y_{n}$ with
@@
where $\mathcal{C}$ is the set of all concave functions $\phi:\mathcal{U}\rightarrow[\underline{B}_{\phi},\bar{B}_{\phi}]$. Note that the problem is not infinite dimensional, indeed we only need to recover $n$ variables $(\phi(U_{1}),\phi(U_{2}),\dots,\phi(U_{n}))\in[\underline{B}_{\phi},\bar{B}_{\phi}]$. When $h(x)=x$ (or equivalently for $1$-concave functions, i.e. concave functions), (41) was studied by Dümbgen et al. (2004), for which $\widehat{\phi}_{n}$ exists and is unique and doesn’t need restrictions on the codomain of $\phi_{0}$, that is $\underline{B}_{\phi},\bar{B}_{\phi}]$ can be the whole $\mathbb{R}$. Existence and uniqueness come from the coercivity of the loss $\mathcal{L}$, which is a property that only applies for $h(x)=x$. Since in our case $h$ is not necessarily the identity map, we rely on Weierstrass’ theorem (which guarantees the existence of at least a minimum for a continuous function $F:\Omega\subset\mathbb{R}^{n}\to\mathbb{R}$ defined on a compact set $\Omega$ of $\mathbb{R}^{n}$). Thus a solution of (41) exists (but might not be unique) because the $n$-dimensional variable $(\phi(U_{1}),\phi(U_{2}),\dots,\phi(U_{n}))\in[\underline{B}_{\phi},\bar{B}_{\phi}]$ where $\underline{B}_{\phi},\bar{B}_{\phi}]$ is compact. Uniqueness is not an issue in our case since we only care about the construction of a consistent estimator. Proposition˜H.1, which is followed by Lemma˜H.7, shows the constraints on the codomain of $\phi_{0}$ when $\psi_{0}:\mathcal{U}\to[\underline{B}_{\psi},\bar{B}_{\psi}]$ is $s$-concave for $s\neq 1$.
@@
We now establish the assumptions to guarantee the convergence of the estimator defined in Equation˜41. We consider a triangular scheme of observations $U_{i}=U_{n,i}$ and $Y_{i}=Y_{n,i}$ but suppress the additional subscript $n$ for notational simplicity. Let $\mathscr{M}_{n}$ be the empirical distribution of the design points $U_{i}$, i.e.
@@
˜H.2 holds if the density $f_{U}$ of the measure $\mathscr{M}$ is bounded away from zero, that is $f_{U}\geq C_{\mathscr{M}}$, which is a standard assumption for uniform convergence of non-parametric functions (see e.g. Dümbgen et al. (2004); Mösching & Dümbgen (2020)) as well as for $L^{2}$ convergence (see e.g. Groeneboom & Hendrickx (2018); Balabdaoui et al. (2019)).
@@
Suppose that ˜H.2, ˜H.3 and ˜H.4 are satisfied. Let $\widehat{\phi}_{n}$ be a solution of (41), and let $\widehat{\psi}_{n}=h\circ\widehat{\phi}_{n}$. Then, for all $\gamma>4$ there exists a integer $n_{0}$ and a real value $\mathscr{C}>0$ such that
@@
The proof can be found in Section˜H.3.
@@
If $\psi_{0}$ and $\widehat{\psi}_{n}$ depends also on a parameter $\bm{\theta}\in\Theta$, $\psi_{0,\bm{\theta}}$ and $\widehat{\psi}_{n,\bm{\theta}}$, for which the domain $\mathcal{U}$ does not depend on $\bm{\theta}$ and such that ˜H.2 and ˜H.4 hold uniformly in $\bm{\theta}\in\Theta$ and all the constants $C_{\mathscr{M}},\bar{B}_{h},\underline{B}_{h},L,\alpha,\gamma,a,b$ are independent of $\bm{\theta}$, then
@@ H.2 Technical Lemma 2
˜H.2 implies the following Lemma˜H.8 that will be used to prove the uniform convergence consistency in Theorem˜H.5.
@@
Let $\rho_{n}\triangleq\frac{\log n}{n}$ and $\text{Leb}(\cdot)$ stands for Lebesgue measure and $U_{1},U_{2},\dots,U_{n}$ i.i.d. points with probability measure $M$ that satisfies ˜H.2, then for a given constant $C_{3}>0$, and for any $\gamma>2$ there exists $n_{0}\in\mathbb{N}$ and a sequence $\epsilon_{n}=\epsilon_{n}(\gamma,C_{3},\alpha)>0$, $\epsilon_{n}\rightarrow 0$ such that
@@
This result immediately follos from the proof of the more general result by Mösching & Dümbgen (2020, Section 4.3) which can be stated as follows:
@@
The next Lemma˜H.10 is proved in Section˜H.4
@@
The next ingredient for our proof is a result concerning the difference of two concave functions, which can be found in Dümbgen et al. (2004, Lemma 5.2) and we report the proof in Section˜H.5 for completeness.
@@
Let $\phi$ satisfy ˜H.4. There is a universal constant $\widetilde{\mathscr{C}}=\widetilde{\mathscr{C}}(L,\alpha)$ with the following property: for any $\xi>0$, let $\delta\leq\widetilde{\mathscr{C}}\min\left(b-a,\xi^{1/\alpha}\right)$. Then, for any concave function $\hat{\phi}$
@@
Now, we have to show that one of our classes $\mathcal{D}_{m}$ does indeed contain useful perturbation functions $\Delta$. We denote with $\check{\phi}$ the unique continuous and piecewise linear function with knots in $\mathscr{T}\cap\left(U_{1},U_{n}\right)$ such that $\check{\phi}_{n}=\widehat{\phi}_{n}$ on $\mathscr{T}=\{U_{1},U_{2},\dots,U_{n}\}$. Thus $\check{\phi}_{n}$ is one particular LS estimator for $\phi_{0}$. The following technical Lemma˜H.12 can be found in Dümbgen et al. (2004, Lemma 3).
@@
From Lemma˜H.11, replacing
@@
Because $c_{n},c_{n}+\delta_{n}]$ is a sub-interval of $\mathcal{U}$ on length $\delta_{n}$, Lemma˜H.8 (which holds by ˜H.2) guarantees that this interval contains at least one observation $U_{i}$. This implies Lemma˜H.12. For any $\Delta\in\mathcal{D}_{6}$ define $I_{\Delta}=\{i:\Delta(U_{i})<0\}$. Define also $J=\{i:\phi_{0}(U_{i})-\widehat{\phi}_{n}(U_{i})>0\}$. Using that $0<h^{\prime}\in[\underline{B}_{h},\bar{B}_{h}]$ with $\underline{B}_{h},\bar{B}_{h}>0$ (which comes from Lemma˜H.7) we have that
@@
Recall that $o_{h}=(\underline{B}_{h},\bar{B}_{h})^{\top}$. By the definition of $S_{n}(6)$ and Lemma˜H.12, there is a (random) function $\Delta\in\mathcal{D}_{6}$ such that
@@
Now consider the event $B_{6,n,m}$ in Lemma˜H.10 and the event $A_{n,\gamma}$ in Lemma˜H.8. For $\gamma>4$ define
@@
then with probability less than $2n^{2-\gamma}$ we have $\sup_{\left[a+\delta_{n},b-\delta_{n}\right]}|\psi_{0}-\widehat{\psi}_{n}|\geq C_{\mathscr{C}}\gamma\delta_{n}^{\alpha}$ provided (by Lemma˜H.10)
@@
More precisely, the condition on $\mathscr{C}$ is that $\mathscr{C}\geq\max\left\{\widetilde{\mathscr{C}}^{-\alpha}\bar{B}_{h},\frac{\gamma\sigma\sqrt{7}}{(4\bar{B}_{h})^{-1}\|o_{h}\|^{-1}\underline{B}_{h}^{2}\left(C_{\mathscr{M}}/8\right)^{1/2}}\right\}$, where $\gamma>4$. Using that $o_{h}=(\underline{B}_{h},\bar{B}_{h})^{\top}$ and $\widetilde{\mathscr{C}}^{\alpha}=\frac{\alpha}{4^{1+\alpha}L}$ (from Lemma˜H.11), this condition becomes
@@
H.4 Proof of Lemma˜H.10
˜H.3 implies the following inequality
@@
Define the concave function $\Phi:=\widehat{\phi}-\widetilde{\phi}$. Now, note that by ˜H.4 we have
@@
    Now suppose that $(\phi-\widehat{\phi})\left(t_{o}\right)\geq\xi$ for some $\xi>0$ and $t_{o}\in[a+\delta,b-\delta]$, where $0<\delta\leq(b-a)/2$. By ˜H.4 and concavity of both $\phi,\widehat{\phi}$ there exist numbers $\gamma,\widehat{\gamma}$ such that
@@ Appendix I Remarks
I.1 Remark on ˜3.2
In this section, we discuss sufficient conditions under which ˜3.2 holds. Under twice differentiability of $\psi_{i}$, the strong concavity of $\operatorname{rev}_{i}\left(\cdot\mid\mathbf{p}_{-i}\right)$ uniformly over $\mathbf{p}_{-i}$, is guaranteed if there exists a value $\mu_{i}>0$, such that $\partial^{2}_{p^{2}}\operatorname{rev}_{i}(p\mid\mathbf{p}_{-i})\geq\mu_{i}$ for every $p\in\mathcal{P}_{i}$, uniformly in $\mathbf{p}_{-i}\in\mathcal{P}_{-i}$. Uniformity in $\mathbf{p}_{-i}\in\mathcal{P}_{-i}$ means that $\mu_{i}$ does not depends on $\mathbf{p}_{-i}$. Here we give some examples where this condition is satisfied.
@@
    Linear demand models (Li et al., 2024). In this case $\operatorname{rev}_{i}(p\mid\mathbf{p}_{-i})=p(\alpha_{i}-\beta_{i}p+\bm{\gamma}_{i}^{\top}\mathbf{p}_{-i})$, and $\partial^{2}_{p^{2}}\operatorname{rev}_{i}(p\mid\mathbf{p}_{-i})=2\beta_{i}>0$.
@@
    Concave demand models. If $\psi_{i}$ is such that $\psi_{i}^{\prime}\geq B_{\psi_{i}^{\prime}}$ for some $B_{\psi_{i}^{\prime}}>0$ as in ˜3.1 and is concave, which implies $\partial_{p}\operatorname{rev}_{i}(p\mid\mathbf{p}_{-i})=\psi_{i}(-\beta_{i}p+\bm{\gamma}_{i}^{\top}\mathbf{p}_{-i})-\beta_{i}p\psi_{i}^{\prime}(-\beta_{i}p+\bm{\gamma}_{i}^{\top}\mathbf{p}_{-i})$, then
@@
    Note that requiring concavity of $\psi_{i}$ is equivalent to requiring 11-concavity of $\psi_{i}$, then, by the inclusion property of $s$-concave functions, any (positive and strictly increasing) $s_{i}$-concave $\psi_{i}$ with $s_{i}\geq 1$ satisfies ˜3.2.
@@
    s-concave demand functions. Suppose that $\psi_{i}$ satisfies ˜3.1 and ˜3.4. Then, by Proposition˜3.5, $\psi_{i}$ is $s_{i}$-concave for some $s_{i}>-1$, and by Lemma˜B.1 this is equivalent to $\psi_{i}^{\prime\prime}\leq(1-s_{i})(\psi_{i}^{\prime})^{2}/\psi_{i}$. This implies that
@@
I.2 Remark˜4.1 on the common exploration phase
To highlight the importance of a common exploration phase, we demonstrate that no seller has an incentive to extend or shorten the exploration phase. If firm $i$ prolongs its exploration phase, it risks incomplete learning of its model parameters (Keskin & Zeevi, 2018), leading to inaccurate parameter estimation and, consequently, a loss in revenue. Conversely, if firm $i$ shortens its exploration phase, it may achieve consistent estimation but at the cost of higher regret due to insufficient exploratory data. For a better illustration, consider a scenario where firm $i$ extends its exploration phase to $\tau^{\prime}>\tau$ while all other firms use a phase of length $\tau$. When firm $i$ estimates its parameters using data $\{(\mathbf{p}^{(t)},y_{i}^{(t)})\}_{t\leq\tau^{\prime}}$, the data from times $\tau+1$ to $\tau^{\prime}$ are no longer iid. In this later period, all other firms have already started their exploration, causing the prices $\{\mathbf{p}_{-i}^{(t)}\}_{t=\tau+1,\dots,\tau^{\prime}}$ to be dependent on the earlier data $\{\mathbf{p}_{-i}^{(t)}\}_{t\leq\tau}$. This dependency can result in inconsistent parameter estimation and a loss of efficiency for firm $i$. Conversely, if firm $i$ opts for a shorter exploration phase than $\tau$, it may achieve consistent estimation, but the reduced amount of exploratory data could lead to lower efficiency compared to firms that adhere to the full exploration phase.
I.3 Remark˜4.2 on the need for two different phases for model estimation
In principle, one could estimate $(\bm{\theta}_{i},\psi_{i})$ jointly using the full exploration phase. For instance, Balabdaoui et al. (2019) employ a profile least squares approach to achieve $L^{2}$ convergence for the joint LSE of $(\bm{\theta}_{i},\psi_{i})$ based on the data $\{(y_{i}^{(t)},\mathbf{p}^{(t)})\}_{t\leq\tau}$. However, in our context, $L^{2}$ convergence alone does not suffice to establish an upper bound on the total expected regret. To elaborate, let $(\bm{\theta}_{i}^{JLS},\psi_{i}^{JLS})$ denote the joint estimator proposed in Balabdaoui et al. (2019) based on the data $\{(y_{i}^{(t)},\mathbf{p}^{(t)})\}_{t\leq\tau}$. Their result applied to our setting gives the following $L^{2}$ convergence rate:
@@
We evaluate the performance of Algorithm˜1 in the sequential price competition model with $N\in\{2,4,6\}$ sellers. Each seller $i\in\mathcal{N}=\{1,2,\dots,N\}$ sets prices supported on $\mathcal{P}=[0,3]^{\times N}$. We examine how convergence behavior changes as the contraction constant $L_{\bm{\Gamma}}$ varies –specifically, when it is close to 0, near $0.5$, or approaches $1$.
@@
which simplifies to $L_{\bm{\Gamma}}=\sup_{i\in\mathcal{N}}\frac{\left\|\gamma_{i}\right\|_{1}}{\beta_{i}}$, when the link functions $\psi_{i}$ are log-concave (see Remark˜3.8). In our experiments, we take
@@
A summary of the two extreme cases can be found in Table˜3.
@@
Given the discussion above, and the table in Table˜3, for each $N\in\{2,4,6\}$, we run simulations under three representative values of $\beta_{i}$, yielding a total of $9$ settings:
@@
In all experiments, demand noise follows $\text{Unif}[-0.05,0.05]$. The common exploration phase has length $\frac{2}{3}T^{5/7}$ and the length $\kappa_{i}$ of the first part of the exploration phase is determined using Equation˜14 with $\mathscr{C}_{i}=1$, $\mathscr{B}_{i}=1$. During the first phase of the exploration phase, the values of $\mathbf{p}^{(t)}$ are samples from a multivariate Gaussian with mean
@@
For each horizon $T\in\{100,400,800,1600,3200\}$, we apply Algorithm˜1 and repeat the simulation $30$ times to obtain averages and $95\%$ confidence intervals.
The results are displayed in Figure˜6, and a log-log version to highlight the convergence rate in Figure˜6. As expected, convergence accelerates as the contraction constant approaches 11, in line with the intuition highlighted in (9).
J.1 Additional Simulation: the different exploration phases across sellers
In this experiment, we replicate the setup of the previous section but relax the common exploration phase assumption in order to test the robustness of our algorithm. An illustration of misaligned exploration phases is provided in Figure˜4.
Figure 4: Illustration of our policy (Algorithm˜1) with $N=4$ sellers in sequential price competition under nonlinear demands. For each seller $i\in\{1,2,3,4\}$, in their exploration phase (dotted line) of length $\tau_{i}$, they offer randomized prices following their distribution $\mathscr{D}_{i}$. Within the exploration phase, each seller has a private phase for estimating $\bm{\theta}_{i}$ (blue box with dotted border), with length $\tau_{i}\kappa_{i}$ and a private phase for estimating $\psi_{i}$ (yellow box with continued border line), with length $\tau_{i}(1-\kappa_{i})$. The $i$-th seller’s price experiment ends at period $t=\tau_{i}$ (black circles). Subsequently, in their exploitation phase (represented by the continued black line), seller $i$ offers prices based on the estimators generated in the exploration phase.

### v4

@@ Appendix H NPLS for $s$-concave regression
Inspired by the work of dumbgen2004consistency, we prove a uniform convergence result for the NPLSE of a mean function that can be written as an increasing function $h$ composed with a concave function $\phi_{0}$. Suppose that we observe $\left(U_{1},Y_{1}\right),\left(U_{2},Y_{2}\right),\ldots,\left(U_{n},Y_{n}\right)$ with fixed numbers $U_{1}\leq U_{2}\leq$ $\cdots\leq U_{n}$ and independent random variables $Y_{1},Y_{2},\ldots,Y_{n}$ with
@@
where $\mathcal{C}$ is the set of all concave functions $\phi:\mathcal{U}\rightarrow[\underline{B}_{\phi},\bar{B}_{\phi}]$. Note that the problem is not infinite dimensional, indeed we only need to recover $n$ variables $(\phi(U_{1}),\phi(U_{2}),\dots,\phi(U_{n}))\in[\underline{B}_{\phi},\bar{B}_{\phi}]$. When $h(x)=x$ (or equivalently for $1$-concave functions, i.e. concave functions), (41) was studied by dumbgen2004consistency, for which $\widehat{\phi}_{n}$ exists and is unique and doesn’t need restrictions on the codomain of $\phi_{0}$, that is $\underline{B}_{\phi},\bar{B}_{\phi}]$ can be the whole $\mathbb{R}$. Existence and uniqueness come from the coercivity of the loss $\mathcal{L}$, which is a property that only applies for $h(x)=x$. Since in our case $h$ is not necessarily the identity map, we rely on Weierstrass’ theorem (which guarantees the existence of at least a minimum for a continuous function $F:\Omega\subset\mathbb{R}^{n}\to\mathbb{R}$ defined on a compact set $\Omega$ of $\mathbb{R}^{n}$). Thus a solution of (41) exists (but might not be unique) because the $n$-dimensional variable $(\phi(U_{1}),\phi(U_{2}),\dots,\phi(U_{n}))\in[\underline{B}_{\phi},\bar{B}_{\phi}]$ where $\underline{B}_{\phi},\bar{B}_{\phi}]$ is compact. Uniqueness is not an issue in our case since we only care about the construction of a consistent estimator. Proposition H.1, which is followed by Lemma H.7, shows the constraints on the codomain of $\phi_{0}$ when $\psi_{0}:\mathcal{U}\to[\underline{B}_{\psi},\bar{B}_{\psi}]$ is $s$-concave for $s\neq 1$.
@@
We now establish the assumptions to guarantee the convergence of the estimator defined in Equation 41. We consider a triangular scheme of observations $U_{i}=U_{n,i}$ and $Y_{i}=Y_{n,i}$ but suppress the additional subscript $n$ for notational simplicity. Let $\mathscr{M}_{n}$ be the empirical distribution of the design points $U_{i}$, i.e.
@@
H.2 holds if the density $f_{U}$ of the measure $\mathscr{M}$ is bounded away from zero, that is $f_{U}\geq C_{\mathscr{M}}$, which is a standard assumption for uniform convergence of non-parametric functions (see e.g. dumbgen2004consistency; mosching2020monotone) as well as for $L^{2}$ convergence (see e.g. groeneboom2018current; balabdaoui2019least).
@@
Suppose that H.2, H.3 and H.4 are satisfied. Let $\widehat{\phi}_{n}$ be a solution of (41), and let $\widehat{\psi}_{n}=h\circ\widehat{\phi}_{n}$. Then, for all $\gamma>4$ there exists a integer $n_{0}$ and a real value $\mathscr{C}>0$ such that
@@
The proof can be found in Section H.3.
@@
If $\psi_{0}$ and $\widehat{\psi}_{n}$ depends also on a parameter $\bm{\theta}\in\Theta$, $\psi_{0,\bm{\theta}}$ and $\widehat{\psi}_{n,\bm{\theta}}$, for which the domain $\mathcal{U}$ does not depend on $\bm{\theta}$ and such that H.2 and H.4 hold uniformly in $\bm{\theta}\in\Theta$ and all the constants $C_{\mathscr{M}},\bar{B}_{h},\underline{B}_{h},L,\alpha,\gamma,a,b$ are independent of $\bm{\theta}$, then
@@ H.2 Technical Lemma 2
H.2 implies the following Lemma H.8 that will be used to prove the uniform convergence consistency in Theorem H.5.
@@
Let $\rho_{n}\triangleq\frac{\log n}{n}$ and $\text{Leb}(\cdot)$ stands for Lebesgue measure and $U_{1},U_{2},\dots,U_{n}$ i.i.d. points with probability measure $M$ that satisfies H.2, then for a given constant $C_{3}>0$, and for any $\gamma>2$ there exists $n_{0}\in\mathbb{N}$ and a sequence $\epsilon_{n}=\epsilon_{n}(\gamma,C_{3},\alpha)>0$, $\epsilon_{n}\rightarrow 0$ such that
@@
This result immediately follos from the proof of the more general result by mosching2020monotone which can be stated as follows:
@@
The next Lemma H.10 is proved in Section H.4
@@
The next ingredient for our proof is a result concerning the difference of two concave functions, which can be found in dumbgen2004consistency and we report the proof in Section H.5 for completeness.
@@
Let $\phi$ satisfy H.4. There is a universal constant $\widetilde{\mathscr{C}}=\widetilde{\mathscr{C}}(L,\alpha)$ with the following property: for any $\xi>0$, let $\delta\leq\widetilde{\mathscr{C}}\min\left(b-a,\xi^{1/\alpha}\right)$. Then, for any concave function $\hat{\phi}$
@@
Now, we have to show that one of our classes $\mathcal{D}_{m}$ does indeed contain useful perturbation functions $\Delta$. We denote with $\check{\phi}$ the unique continuous and piecewise linear function with knots in $\mathscr{T}\cap\left(U_{1},U_{n}\right)$ such that $\check{\phi}_{n}=\widehat{\phi}_{n}$ on $\mathscr{T}=\{U_{1},U_{2},\dots,U_{n}\}$. Thus $\check{\phi}_{n}$ is one particular LS estimator for $\phi_{0}$. The following technical Lemma H.12 can be found in dumbgen2004consistency.
@@
From Lemma H.11, replacing
@@
Because $c_{n},c_{n}+\delta_{n}]$ is a sub-interval of $\mathcal{U}$ on length $\delta_{n}$, Lemma H.8 (which holds by H.2) guarantees that this interval contains at least one observation $U_{i}$. This implies Lemma H.12. For any $\Delta\in\mathcal{D}_{6}$ define $I_{\Delta}=\{i:\Delta(U_{i})<0\}$. Define also $J=\{i:\phi_{0}(U_{i})-\widehat{\phi}_{n}(U_{i})>0\}$. Using that $0<h^{\prime}\in[\underline{B}_{h},\bar{B}_{h}]$ with $\underline{B}_{h},\bar{B}_{h}>0$ (which comes from Lemma H.7) we have that
@@
Recall that $o_{h}=(\underline{B}_{h},\bar{B}_{h})^{\top}$. By the definition of $S_{n}(6)$ and Lemma H.12, there is a (random) function $\Delta\in\mathcal{D}_{6}$ such that
@@
Now consider the event $B_{6,n,m}$ in Lemma H.10 and the event $A_{n,\gamma}$ in Lemma H.8. For $\gamma>4$ define
@@
then with probability less than $2n^{2-\gamma}$ we have $\sup_{\left[a+\delta_{n},b-\delta_{n}\right]}|\psi_{0}-\widehat{\psi}_{n}|\geq C_{\mathscr{C}}\gamma\delta_{n}^{\alpha}$ provided (by Lemma H.10)
@@
More precisely, the condition on $\mathscr{C}$ is that $\mathscr{C}\geq\max\left\{\widetilde{\mathscr{C}}^{-\alpha}\bar{B}_{h},\frac{\gamma\sigma\sqrt{7}}{(4\bar{B}_{h})^{-1}\|o_{h}\|^{-1}\underline{B}_{h}^{2}\left(C_{\mathscr{M}}/8\right)^{1/2}}\right\}$, where $\gamma>4$. Using that $o_{h}=(\underline{B}_{h},\bar{B}_{h})^{\top}$ and $\widetilde{\mathscr{C}}^{\alpha}=\frac{\alpha}{4^{1+\alpha}L}$ (from Lemma H.11), this condition becomes
@@
H.4 Proof of Lemma H.10
H.3 implies the following inequality
@@
Define the concave function $\Phi:=\widehat{\phi}-\widetilde{\phi}$. Now, note that by H.4 we have
@@
    Now suppose that $(\phi-\widehat{\phi})\left(t_{o}\right)\geq\xi$ for some $\xi>0$ and $t_{o}\in[a+\delta,b-\delta]$, where $0<\delta\leq(b-a)/2$. By H.4 and concavity of both $\phi,\widehat{\phi}$ there exist numbers $\gamma,\widehat{\gamma}$ such that
@@ Appendix I Remarks
I.1 Remark on 3.2
In this section, we discuss sufficient conditions under which 3.2 holds. Under twice differentiability of $\psi_{i}$, the strong concavity of $\operatorname{rev}_{i}\left(\cdot\mid\mathbf{p}_{-i}\right)$ uniformly over $\mathbf{p}_{-i}$, is guaranteed if there exists a value $\mu_{i}>0$, such that $\partial^{2}_{p^{2}}\operatorname{rev}_{i}(p\mid\mathbf{p}_{-i})\geq\mu_{i}$ for every $p\in\mathcal{P}_{i}$, uniformly in $\mathbf{p}_{-i}\in\mathcal{P}_{-i}$. Uniformity in $\mathbf{p}_{-i}\in\mathcal{P}_{-i}$ means that $\mu_{i}$ does not depends on $\mathbf{p}_{-i}$. Here we give some examples where this condition is satisfied.
@@
    Linear demand models (li2024lego). In this case $\operatorname{rev}_{i}(p\mid\mathbf{p}_{-i})=p(\alpha_{i}-\beta_{i}p+\bm{\gamma}_{i}^{\top}\mathbf{p}_{-i})$, and $\partial^{2}_{p^{2}}\operatorname{rev}_{i}(p\mid\mathbf{p}_{-i})=2\beta_{i}>0$.
@@
    Concave demand models. If $\psi_{i}$ is such that $\psi_{i}^{\prime}\geq B_{\psi_{i}^{\prime}}$ for some $B_{\psi_{i}^{\prime}}>0$ as in 3.1 and is concave, which implies $\partial_{p}\operatorname{rev}_{i}(p\mid\mathbf{p}_{-i})=\psi_{i}(-\beta_{i}p+\bm{\gamma}_{i}^{\top}\mathbf{p}_{-i})-\beta_{i}p\psi_{i}^{\prime}(-\beta_{i}p+\bm{\gamma}_{i}^{\top}\mathbf{p}_{-i})$, then
@@
    Note that requiring concavity of $\psi_{i}$ is equivalent to requiring 11-concavity of $\psi_{i}$, then, by the inclusion property of $s$-concave functions, any (positive and strictly increasing) $s_{i}$-concave $\psi_{i}$ with $s_{i}\geq 1$ satisfies 3.2.
@@
    s-concave demand functions. Suppose that $\psi_{i}$ satisfies 3.1 and 3.4. Then, by Proposition 3.5, $\psi_{i}$ is $s_{i}$-concave for some $s_{i}>-1$, and by Lemma B.1 this is equivalent to $\psi_{i}^{\prime\prime}\leq(1-s_{i})(\psi_{i}^{\prime})^{2}/\psi_{i}$. This implies that
@@
I.2 Remark 4.1 on the common exploration phase
To highlight the importance of a common exploration phase, we demonstrate that no seller has an incentive to extend or shorten the exploration phase. If firm $i$ prolongs its exploration phase, it risks incomplete learning of its model parameters (keskin2018incomplete), leading to inaccurate parameter estimation and, consequently, a loss in revenue. Conversely, if firm $i$ shortens its exploration phase, it may achieve consistent estimation but at the cost of higher regret due to insufficient exploratory data. For a better illustration, consider a scenario where firm $i$ extends its exploration phase to $\tau^{\prime}>\tau$ while all other firms use a phase of length $\tau$. When firm $i$ estimates its parameters using data $\{(\mathbf{p}^{(t)},y_{i}^{(t)})\}_{t\leq\tau^{\prime}}$, the data from times $\tau+1$ to $\tau^{\prime}$ are no longer iid. In this later period, all other firms have already started their exploration, causing the prices $\{\mathbf{p}_{-i}^{(t)}\}_{t=\tau+1,\dots,\tau^{\prime}}$ to be dependent on the earlier data $\{\mathbf{p}_{-i}^{(t)}\}_{t\leq\tau}$. This dependency can result in inconsistent parameter estimation and a loss of efficiency for firm $i$. Conversely, if firm $i$ opts for a shorter exploration phase than $\tau$, it may achieve consistent estimation, but the reduced amount of exploratory data could lead to lower efficiency compared to firms that adhere to the full exploration phase.
I.3 Remark 4.2 on the need for two different phases for model estimation
In principle, one could estimate $(\bm{\theta}_{i},\psi_{i})$ jointly using the full exploration phase. For instance, balabdaoui2019least employ a profile least squares approach to achieve $L^{2}$ convergence for the joint LSE of $(\bm{\theta}_{i},\psi_{i})$ based on the data $\{(y_{i}^{(t)},\mathbf{p}^{(t)})\}_{t\leq\tau}$. However, in our context, $L^{2}$ convergence alone does not suffice to establish an upper bound on the total expected regret. To elaborate, let $(\bm{\theta}_{i}^{JLS},\psi_{i}^{JLS})$ denote the joint estimator proposed in balabdaoui2019least based on the data $\{(y_{i}^{(t)},\mathbf{p}^{(t)})\}_{t\leq\tau}$. Their result applied to our setting gives the following $L^{2}$ convergence rate:
@@
J.1 Contraction constant varying regimes
We evaluate the performance of Algorithm 1 in the sequential price competition model with $N\in\{2,4,6\}$ sellers. Each seller $i\in\mathcal{N}=\{1,2,\dots,N\}$ sets prices supported on $\mathcal{P}=[0,3]^{\times N}$. We examine how convergence behavior changes as the contraction constant $L_{\bm{\Gamma}}$ varies –specifically, when it is close to 0, near $0.5$, or approaches $1$.
@@
which simplifies to $L_{\bm{\Gamma}}=\sup_{i\in\mathcal{N}}\frac{\left\|\gamma_{i}\right\|_{1}}{\beta_{i}}$, when the link functions $\psi_{i}$ are log-concave (see Remark 3.8). In our experiments, we take
@@
A summary of the two extreme cases can be found in Table 3.
@@
Given the discussion above, and the table in Table 3, for each $N\in\{2,4,6\}$, we run simulations under three representative values of $\beta_{i}$, yielding a total of $9$ settings:
@@
In all experiments, demand noise follows $\text{Unif}[-0.05,0.05]$. The common exploration phase has length $\frac{2}{3}T^{5/7}$ and the length $\kappa_{i}$ of the first part of the exploration phase is determined using Equation 14 with $\mathscr{C}_{i}=1$, $\mathscr{B}_{i}=1$. During the first phase of the exploration phase, the values of $\mathbf{p}^{(t)}$ are samples from a multivariate Gaussian with mean
@@
For each horizon $T\in\{100,400,800,1600,3200\}$, we apply Algorithm 1 and repeat the simulation $30$ times to obtain averages and $95\%$ confidence intervals.
The results are displayed in Figure 6, and a log-log version to highlight the convergence rate in Figure 6. As expected, convergence accelerates as the contraction constant approaches 11, in line with the intuition highlighted in (9).
J.2 Different exploration phases across sellers
In this experiment, we replicate the setup in Section J.1 but relax the common exploration phase assumption in order to test the robustness of our algorithm. An illustration of misaligned exploration phases is provided in Figure 4.
Figure 4: Illustration of our policy (Algorithm 1) with $N=4$ sellers in sequential price competition under nonlinear demands. For each seller $i\in\{1,2,3,4\}$, in their exploration phase (dotted line) of length $\tau_{i}$, they offer randomized prices following their distribution $\mathscr{D}_{i}$. Within the exploration phase, each seller has a private phase for estimating $\bm{\theta}_{i}$ (blue box with dotted border), with length $\tau_{i}\kappa_{i}$ and a private phase for estimating $\psi_{i}$ (yellow box with continued border line), with length $\tau_{i}(1-\kappa_{i})$. The $i$-th seller’s price experiment ends at period $t=\tau_{i}$ (black circles). Subsequently, in their exploitation phase (represented by the continued black line), seller $i$ offers prices based on the estimators generated in the exploration phase.